OpenStudy (anonymous):

whats the domain of g(x) = sqrt of (7-3x) ?

OpenStudy (anonymous):

negative infinity, 2)

OpenStudy (anonymous):

i know that to make it zero it is 7/3

OpenStudy (anonymous):

yep that sound more like it. \[(-\infty,\frac{7}{3})\]

OpenStudy (anonymous):

\[7-3x\geq 0\] \[7\geq 3x\] \[\frac{7}{3}\geq x\]

OpenStudy (anonymous):

k thanks so much!